# Problem Solved? Information Extraction Design Space for Layout-Rich Documents using LLMs

### v1

@@ Abstract
This paper defines and explores the design space for information extraction (IE) from layout-rich documents using large language models (LLMs). The three core challenges of layout-aware IE with LLMs are 1) data structuring, 2) model engagement, and 3) output refinement. Our study delves into the sub-problems within these core challenges, such as input representation, chunking, prompting, and selection of LLMs and multimodal models. It examines the outcomes of different design choices through a new layout-aware IE test suite, benchmarking against the state-of-art (SoA) model LayoutLMv3. The results show that the configuration from one-factor-at-a-time (OFAT) trial achieves near-optimal results with $14.1$ points F1-score gain from the baseline model, while full factorial exploration yields only a slightly higher $15.1$ points gain at 36$\sim$ 36$\times$ 36 × greater token usage. We demonstrate that well-configured general-purpose LLMs can match the performance of specialized models, providing a cost-effective alternative. Our test-suite is freely available at https://github.com/gayecolakoglu/LayIE-LLM.
@@
Gaye Colakoglu1,2  Gürkan Solmaz2  Jonathan Fürst1 1Zurich University of Applied Sciences, Switzerland 2NEC Laboratories Europe, Heidelberg, Germany colgay01@students.zhaw.ch, gurkan.solmaz@neclab.eu, jonathan.fuerst@zhaw.ch
@@ 1 Introduction
Information extraction (IE) entails extracting structured data, such as names, dates, or financial figures, from unstructured documents. Within organizations, key information often resides in layout-rich documents (LRDs) such as reports and presentations that combine visual elements (e.g., charts, tables) with textual structure and content Park et al. (2019); Wang et al. (2023a); Zmigrod et al. (2024b). LRDs challenge traditional natural language processing (NLP) techniques, which are designed for plain texts Cui et al. (2021); Tang et al. (2023).
@@
Figure 1: Design space for information extraction from layout-rich documents using LLMs. The goal is to extract all relevant information to the target data schema with correct mapping.
@@
Recent layout-aware models at the intersection of NLP and Computer Vision (CV) address this gap by including visual and structural features to improve information extraction from LRDs (e.g., LayoutLMv1-v3 Xu et al. (2020b, a); Huang et al. (2022), GraphDoc Zhang et al. (2022), DocFormer Appalaraju et al. (2021)). However, these models require substantial dataset-specific fine-tuning; users must manually annotate a training dataset with bounding boxes and extraction elements for each new document set.
@@
While LLMs show significant potential, there exist open questions regarding about their adoption for IE from LRDs. In which form should the document content be provided to the LLM? Which methods are most effective for in-context learning (ICL) and instruction tuning? Do larger multimodal LLMs like GPT-4o provide advantages over traditional Optical Character Recognition (OCR) and text-based models? How coherent is the LLM output, and what post-processing is needed? How does the performance of open- and closed-sourced LLMs of different scales compare? How does performance compare to fine-tuned SoA layout-aware models?
This paper addresses the questions above, by systematically exploring the design space for IE with LLMs from LRDs. We implement and evaluate a range of preprocessing, chunking, prompting, and post-processing techniques alongside ICL strategies. Furthermore, we experimentally assess the performance of diverse LLMs and LayoutLMv3 Huang et al. (2022), a fine-tuned layout-aware model.
@@
Our results reveal multiple important insights. First, current, general LLMs can easily compete with SoA fine-tuned models such as LayoutLMv3 with the additional benefit to not require any training data. Second, instead of fine-tuning through data, LLMs require tuning of the IE pipeline to achieve a competitive performance—the gap between a best-practices baseline and the tuned configuration achieved by our lightweight OFAT method is 14.1 points in F1. Third, while purely text-based LLMs achieve a competitive performance through our method, multi-modal LLMs that directly combine textual with visual features still achieve higher performance with the disadvantage of higher costs (token use and API costs) and less transparency.
Shortly, our contributions are as follows:
@@
  • Introducing the Design Space of IE from LRDs using LLMs, consisting of the three core challenges: 1) Data structuring, 2) Model engagement, and 3) Output refinement (Sec. 2).
@@
  • Development of a layout-aware IE test suite for analyzing effects of OCR- and text-based inputs, chunk sizes, few-shots and CoT prompting, LLM model selection, decoding, entity mapping, data cleaning, and F1-score-based benchmarking using exact, substring, and fuzzy matches (Sec. 3).
  • Comprehensive evaluation using GPT-4o, GPT3.5, and LLaMA3 models, GPT4-vision, and the SoA LayoutLMv3 model (Sec. 4).
  • We open-source all our code and experimental results (data accessible in Wang et al. (2023b)). Our test suite can be used and adapted by others to tune their IE pipelines to their datasets and LLMs (https://github.com/gayecolakoglu/LayIE-LLM).
2 Design Space of IE from Layout-rich Documents with LLMs
@@
2.1 Using LLMs for Information Extraction
IE systems that utilize LLMs have to tackle the three main challenge areas that we consider as part of the Design Space: Data Structuring, Model Engagement, and Output Refinement. Each stage, in their respective order, plays an important role for having an IE system with satisfactory accuracy and robustness.
@@ 2.1 Using LLMs for Information Extraction
Data Structuring. For multimodal LLMs, LRDs can be directly given as input. On the other hand, for purely text-based LLM, the input documents must be transformed into textual representations. This involves converting documents into machine-readable formats using OCR systems to extract features such as text, bounding boxes, and visual elements Mieskes and Schmunk (2019); Smith (2007). Alternatively, a formatting language such as Markdown can be employed to represent the document’s layout, allowing the LLM to understand the structural context of the text better. The impact of OCR quality on IE performance has been documented Bhadauria et al. (2024), and structured formats tend to yield better results Bai et al. (2024). To process larger documents efficiently, they are often divided into smaller, manageable chunks based on page boundaries, sections, or semantic units Liu et al. (2024). Markdown as an input format compared to raw OCR outputs remains underexplored, representing a potential research gap in IE system development.
@@
Model Engagement. Once preprocessed, the document is fed to an LLM for IE. Ensuring alignment between the extracted text and layout information is crucial for accurate representation Xu et al. (2020a); Appalaraju et al. (2021). Prompt-driven extraction leverages general-purpose models, using tailored prompts to guide the extraction process Brown et al. (2020); Radford et al. (2021); Zhou et al. (2022). As such, the model needs to be instructed to extract information, and usually, the schema is provided. This step can involve more advanced IT and ICL techniques (few-shot, CoT). The influence of prompting techniques in interaction with various stages of the IE pipeline to enhance performance and robustness remains a research gap that requires further investigation.
@@
Output Refinement. After inference, the extracted information undergoes post-processing to ensure accuracy and conformance to Schema S. This step involves refining and validating the outputs generated by the LLM through tasks such as mapping extracted entities E to their original document positions, merging overlapping or fragmented predictions, and resolving ambiguities in the results Xu et al. (2020b). Refining entity extraction through post-processing has been explored in various studies Wang et al. (2022b); Tamayo et al. (2022). Rule-based entity alignment has notable accuracy improvements Luo et al. (2024). There exists no analysis of post-processing techniques tailored to LRDs in conjunction with LLMs.
3 Test-Suite for IE from Layout-rich Documents with LLMs
@@
We implement a comprehensive test suite to assess IE tasks from LRDs. The pipeline, depicted in Figure 2, systematically transforms raw input into structured output across multiple stages for evaluating the efficacy of design decisions.
@@
Figure 2: Overview of the test suite for extracting information from layout-rich documents using LLMs in six stages. The process begins with OCR-based text extraction and Markdown conversion with LLM assistance, followed by chunking to manage token limits, experimenting with different chunk sizes. Each chunk is processed to generate a prompt, exploring Few-shot and CoT prompting with varying numbers of examples. These examples consist of a document example and extracted key-value pairs in a structured format, along with a new document and the required task for the LLM. LLMs generate structured JSON outputs, which are then decoded and reconciled. Post-processing involves experimenting data cleaning and entity mapping. Finally, evaluation is conducted using two different evaluation techniques.
@@ 3.1 Data Structuring
We convert raw documents into machine-readable formats by extracting text and layout features. The process involves two conversions: 1) Extracting textual content and layout information using OCR data, and 2) creating a markdown representation of PDFs.
Chunking. We employ three chunk sizes: (1) max: 4096 tokens, (2) medium: 2048 tokens, and (3) small: 1024 tokens. Documents are segmented into $N$ chunks based on document length and chunk size, preparing them for the prompting phase. Each chunk is formed by accumulating whole words until the token limit. This ensures intact word boundaries and maintaining a sequential, non-overlapping structure. Layout information is preserved by associating each text segment with normalized and quantized spatial coordinates, retaining the structural context of the document.
@@ 3.2 Model Engagement
Model engagement consists of constructing input to the LLM comprising at least three components: (1) a task instruction outlining the IE task, (2) the target schema $S$, and (3) the document chunk. We adhere to best practices from NLP for prompt structure and IE task instruction. The schema $S$ is implemented as a dictionary of key-value pairs, where values specify the format of the corresponding attribute using regex expressions in Listing 1.
@@ 3.3 Output Refinement
We refine the raw outputs from the LLM to ensure alignment with the target schema, addressing challenges related to prediction variability, schema definition differences, and data formatting inconsistencies (e.g., varying date formats). We implement three techniques inspired by related work in data integration Dong and Srivastava (2013): Decoding, Schema Mapping, and Data Cleaning. This process results in three sets of predictions: initial predictions, mapped predictions, and cleaned predictions.
Decoding. The decoding step parses each LLM completion as a JSON object, discarding any that fail to parse. The process then consolidates predictions for each document by reconciling outputs generated across individual pages and chunks. With $N$ completions for $N$ prompts, corresponding to $N$ chunks, the model generates multiple predictions for a single document. Reconciliation ensures a unified document-level output by deduplicating nested predictions and aggregating unique values. If multiple unique values exist for a single entity, they are stored together to preserve variability. The outcome of this step is referred to as the initial predictions.
@@
  • Fuzzy Match uses similarity metrics for approximate matches. A match is valid if the highest similarity ratio exceeds a predefined threshold (default: 0.8). This method is well-suited for scenarios with minor variations caused by OCR errors or formatting discrepancies.
@@ 4.1 Experimental Design
Methodology.
The goal of our experimental setup is to study how different parameters in the pipeline, shown in Figure 2, affect the overall performance of IE from LRDs using LLMs. To investigate the design dimensions, we start with a baseline configuration and and systematically alter factors at a single dimension at a time, following a one-factor-at-a-time (OFAT) methodology. This approach allows us to isolate and understand the impact of each parameter change on the IE performance.
Our intuition is that aggregating the knowledge gained for each dimension independently, we can achieve a deeper understanding of the design space and possibly identify an effective overall configuration for IE, without the need for a comprehensive factorial exploration. However, we also validate the findings by comparing the results of the OFAT method with those obtained from a brute-force approach, which is based on conducting 432 experiments.
Dataset and LLMs.
We utilize the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023b), which includes two benchmarks. Each benchmark includes training samples of 10, 50, 100, and 200 documents with high-quality OCR for assessing data efficiency, as well as generalization tasks: Single Template Learning (STL), Unseen Template Learning (UTL), and Mixed Template Learning (MTL). For further details on how this dataset is tailored for our experiments and diverse models, please refer to Appendix A.2.
We evaluate GPT-3.5, GPT-4o, and LLaMA3-70B for text-only structured data extraction from LRDs. Additionally, we compare their results with GPT-4 Vision and LayoutLMv3 to assess the performance gap between multimodal LLMs and domain-specific, fine-tuned models, respectively.
The Baseline Configuration.
The baseline configuration is outlined in Table 1,where the configuration is selected based on best practices such as in Perot et al. (2024) for the following reasons: (1) OCR reflects real-world scenarios for digitized LRDs. (2) Medium chunk size balances efficiency and context preservation, addressing token limits in LLMs. (3) Few-shot prompting combines pre-trained knowledge with minimal task-specific guidance. (4) Using zero examples provides a clear benchmark for assessing the model’s raw performance. (5) Initial predictions are retained to evaluate models’ raw output without modifications, ensuring a direct assessment of their capabilities. (6) Finally, exact match provides a stringent measure of correctness, offering a reliable baseline for comparison across configurations.

### v2

@@ Abstract
This paper defines and explores the design space for information extraction (IE) from layout-rich documents using large language models (LLMs). The three core challenges of layout-aware IE with LLMs are 1) data structuring, 2) model engagement, and 3) output refinement. Our study investigates the sub-problems and methods within these core challenges, such as input representation, chunking, prompting, selection of LLMs, and multimodal models. It examines the effect of different design choices through LayIE-LLM, a new, open-source, layout-aware IE test suite, benchmarking against traditional, fine-tuned IE models. The results on two IE datasets show that LLMs require adjustment of the IE pipeline to achieve competitive performance: the optimized configuration found with LayIE-LLM achieves 13.3–37.5 F1 points more than a general-practice baseline configuration using the same LLM. To find a well-working configuration, we develop a one-factor-at-a-time (OFAT) method that achieves near-optimal results. Our method is only 0.8–1.8 points lower than the best full factorial exploration with a fraction ($\sim$2.8%) of the required computation. Overall, we demonstrate that, if well-configured, general-purpose LLMs match the performance of specialized models, providing a cost-effective, finetuning-free alternative. Our test-suite is available at https://github.com/gayecolakoglu/LayIE-LLM.
@@
Gaye Colakoglu1,2  Gürkan Solmaz2  Jonathan Fürst1 1Zurich University of Applied Sciences, Switzerland 2NEC Laboratories Europe, Heidelberg, Germany colgay01@students.zhaw.ch, guerkan.solmaz@neclab.eu, jonathan.fuerst@zhaw.ch
@@ 1 Introduction
Information extraction (IE) extracts structured data from unstructured documents, including layout-rich documents (LRDs) as reports and presentations that mix visual and textual elements Park et al. (2019); Wang et al. (2023b); Zmigrod et al. (2024b). LRDs challenge traditional natural language processing (NLP) techniques, designed for plain texts Cui et al. (2021); Tang et al. (2023).
@@
Figure 1: Design space for IE from layout-rich documents using LLMs. The goal is to extract information relevant to the target data schema with correct mapping.
@@
Recent layout-aware models at the intersection of NLP and Computer Vision (CV) address this gap by including visual and structural features to improve IE from LRDs (e.g., LayoutLMv1-v3 Xu et al. (2020b, a); Huang et al. (2022), ERNIE-Layout Peng et al. (2022), GraphDoc Zhang et al. (2022), DocFormer Appalaraju et al. (2021)). While early layout-aware models required substantial dataset-specific fine-tuning, recent generative document understanding models Powalski et al. (2021) have demonstrated strong zero- and few-shot capabilities on unseen datasets and tasks.
@@
While LLMs exhibit significant potential, there exist open questions regarding their adoption for IE from LRDs. In what form should the document content be provided to the LLM? Which methods are most effective for in-context learning (ICL) and instruction tuning? Do large multimodal LLMs as GPT-4o OpenAI (2024) and Qwen 2.5-vision Bai et al. (2025) offer advantages over traditional Optical Character Recognition (OCR) and fine-tuned layout-aware models, and how does their performance vary across LLMs of different scales and sources? How coherent is the LLM output, and what post-processing is needed?
Studies have individually addressed some of the aspects Wang et al. (2023a); Luo et al. (2024a); Blau et al. (2024). However, there is a lack of a unified framework that comprehensively explores and compares these components. This paper fills this gap by systematically investigating the design space for LLM-based IE from LRDs. We focus on evaluating a range of preprocessing, chunking, prompting, and post-processing techniques, alongside various ICL strategies, within a unified, reproducible framework. Additionally, we conduct an extensive comparison of these components, examining multimodal LLMs, traditional models, open- and closed-source LLMs, and fine-tuned layout-aware models. To the best of our knowledge, no existing research has comprehensively compared the combined impact of these pipeline parameters across such diverse model categories.
@@
Our results reveal multiple insights. First, current general LLMs can compete with traditional fine-tuned models such as LayoutLMv3 and ERNIE-Layout, without expensive labeling. Second, rather than fine-tuning on data, LLMs require adjustment of the IE pipeline to achieve competitive performance–the performance gap between a general-practice baseline and the configuration tuned using our lightweight OFAT method on two datasets is 13.3 F1 points for VRDU Wang et al. (2023c), and 37.5 for FUNSD Jaume et al. (2019), only 1.8 points and 0.8 points lower than the best possible Brute-Force configuration respectively. Third, while purely text-based LLMs achieve competitive performance with our method, multimodal LLMs, those that directly integrate textual and visual features, still outperform them. However, this advantage comes with increased token usage, higher API costs, and diminished transparency and control over individual steps. In summary, our contributions are as follows.
@@
  • We introduce the Design Space of IE from LRDs using LLMs, addressing three core challenges: 1) Data structuring, 2) Model engagement, and 3) Output refinement (Sec. 2).
@@
  • We develop LayIE-LLM, a layout-aware IE test suite to analyze the impact of: OCR, text-based inputs, chunk size, few-shot and CoT prompting, LLM models, decoding strategies, schema mapping, data cleaning, and evaluation techniques using exact, substring, and fuzzy matching (Sec. 3).
  • We conduct evaluation of GPT-4o, GPT-3.5, and LLaMA3, as well as the vision-enabled GPT-4o and Qwen2.5-vision, and compare them with traditional, fine-tuned layout-aware models such as LayoutLMv3 and ERNIE-Layout, highlighting trade-offs between performance and resource efficiency (Sec. 4).
  • We open-source LayIE-LLM and all our experimental results, enabling reproducibility and further experimentation by the community: https://github.com/gayecolakoglu/LayIE-LLM
2 Design Space of IE from LRDs with LLMs
@@
2.1 Using LLMs for IE
LLM-based IE systems have to tackle three challenge areas: Data Structuring, Model Engagement, and Output Refinement.
@@ 2.1 Using LLMs for IE
Data Structuring. For multimodal LLMs, LRDs can be directly given as input. On the other hand, for purely text-based LLMs, the input documents must be transformed into textual representations. This involves converting documents into machine-readable formats using OCR systems to extract features such as text, bounding boxes, and visual elements Mieskes and Schmunk (2019); Smith (2007). Alternatively, a formatting language such as Markdown can be employed to represent the document’s layout, allowing the LLM to understand the structural context of the text better. The impact of OCR quality on IE performance has been documented Bhadauria et al. (2024), and structured formats tend to yield better results Bai et al. (2024). To process larger documents efficiently, they are often divided into smaller, manageable chunks based on page boundaries, sections, or semantic units Liu et al. (2024). Markdown as an input format compared to raw OCR outputs remains underexplored, representing a potential research gap in IE system development.
@@
Model Engagement. Once preprocessed, the document is fed to an LLM for IE. Ensuring alignment between the extracted text and layout information is crucial for accurate representation Xu et al. (2020a); Appalaraju et al. (2021). Prompt-driven extraction leverages general-purpose models, by using tailored prompts to guide the extraction process Brown et al. (2020); Radford et al. (2021); Zhou et al. (2022). As such, models must be explicitly instructed to extract information, typically with an accompanying schema. This step can involve more advanced IT and ICL techniques (few-shot, CoT). The influence of prompting techniques in interaction with various stages of the IE pipeline to enhance performance and robustness remains a research gap that requires further investigation.
@@
Output Refinement. After inference, the extracted information undergoes post-processing to ensure accuracy and conformity with Schema $S$. This step involves refining and validating the outputs generated by the LLM through tasks such as mapping extracted entities $E$ to their original document positions, merging overlapping or fragmented predictions, and resolving ambiguities in the results Xu et al. (2020b). Post-processing for entity extraction has been explored in prior studies Wang et al. (2022b); Tamayo et al. (2022), and rule-based entity alignment has demonstrated notable improvements in accuracy Luo et al. (2024b). However, to our knowledge, there has been no analysis of post-processing techniques specifically tailored to LRDs in conjunction with LLMs.
3 LayIE-LLM: Test-Suite for IE from LRDs with LLMs
@@
We implement LayIE-LLM , a comprehensive test suite to evaluate IE tasks from LRDs. The pipeline, depicted in Figure 2, systematically transforms raw inputs into structured outputs through multiple stages, enabling a thorough assessment of the effectiveness of various design decisions.
@@
Figure 2: LayIE-LLM test suite for extracting information from LRDs using LLMs in six stages. The process begins with OCR-based text extraction and Markdown conversion with LLM assistance, followed by chunking to manage token limits, experimenting with different chunk sizes. Each chunk is processed into a prompt using Few-shot and CoT strategies with varying example counts. Prompts include a document example with key-value pairs, plus a new document and task. LLMs generate structured JSON outputs, which are decoded and reconciled. Post-processing includes data cleaning and entity mapping, followed by evaluation using two methods.
@@ 3.1 Data Structuring
PDF documents include both textual content and layout features. The process involves two conversions: 1) Extracting textual content and layout information using OCR data, and 2) creating a markdown representation.
Chunking. We employ three chunk sizes: (1) max: 4096 tokens, (2) medium: 2048 tokens, and (3) small: 1024 tokens. Documents are segmented into $N$ chunks based on their length and the selected chunk size. Each chunk is constructed by sequentially accumulating whole words up to the token limit, thereby preserving word boundaries and ensuring a non-overlapping structure. Layout information is retained by associating each text segment with normalized and quantized spatial coordinates, which preserve the structural context of the original document.
@@ 3.2 Model Engagement
Model engagement involves constructing input to the LLM using three primary components: (1) a task instruction that defines the IE task, (2) the target schema $S$, and (3) the document chunk. We adhere to best practices from NLP for prompt structure and IE task instruction. The schema $S$ is implemented as a dictionary of key-value pairs, where values specify the format of the corresponding attribute using regex expressions in Listing 1.
@@ 3.3 Output Refinement
We refine the raw outputs from the LLM to ensure alignment with the target schema, addressing challenges related to prediction variability, schema definition differences, and data formatting inconsistencies (e.g., varying date formats). We implement three techniques inspired by related work in data integration Dong and Srivastava (2013): Decoding, Schema Mapping, and Data Cleaning. This results in three sets of predictions: initial predictions, mapped predictions, and cleaned predictions.
Decoding. The decoding step parses each LLM completion as a JSON object, discarding any that fail to parse. The process then consolidates predictions for each document by reconciling outputs generated across individual pages and chunks. With $N$ completions for $N$ prompts, corresponding to $N$ chunks, the model generates multiple predictions for a single document. Reconciliation ensures a unified document-level output by deduplicating nested predictions and aggregating unique values. If multiple unique values exist for a single entity, they are retained to preserve variability. The outcome of this step is referred to as initial predictions.
@@
  • Fuzzy Match uses similarity metrics (fuzz.ratio) for approximate matches. A match is valid if the highest similarity ratio exceeds a predefined threshold (default: 0.8), creating binary decisions unlike soft similarity metrics Biten et al. (2019). This method is well-suited for scenarios with minor variations caused by OCR errors or formatting discrepancies.
@@ 4.1 Experimental Design
Methodology. The goal of our experimental setup is to analyze how different parameters in the pipeline (Figure 2) influence the performance of IE from LRDs using LLMs. To examine the design dimensions, we begin with a baseline configuration and systematically vary one factor at a time, following a one-factor-at-a-time (OFAT) methodology. This approach allows us to isolate and understand the impact of each parameter change on the IE performance.
Our intuition is that by independently aggregating the insights gained from each design dimension, one can develop a deeper understanding of the design space and potentially identify an effective overall configuration for IE, without the need for an exhaustive factorial exploration. To validate this approach, we compare the OFAT method’s results with those obtained via a brute-force strategy involving 432 experiments ($2\times 3\times 2\times 4\times 3\times 3=432$, see Table 1).
Dataset and LLMs. We use the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023c), which includes two benchmarks with high-quality OCR for assessing data efficiency, due to its inclusion of three visually diverse template types – Amendment, Short-Form, and Dissemination – representing varied and complex layouts, as well as generalization tasks: Single Template Learning (STL), Unseen Template Learning (UTL), and Mixed Template Learning (MTL). For further details on how this dataset is tailored for our experiments and diverse models, please refer to A.3.
Additionally, to further prove the generalizability of our approach, we also conducted the same set of experiments using the FUNSD dataset Jaume et al. (2019), which directly provides key-value pairs in a question-answering format without the complexity of OCR processing. Due to page limitations and for the sake of clarity, the experimental results presented in this section are based on the VRDU dataset. The corresponding experimental results for the FUNSD dataset are provided in A.5.
We evaluate GPT-3.5, GPT-4o, and LLaMA3-70B for text-only structured data extraction from LRDs. Furthermore, we compare their results with those of GPT-4, LayoutLMv3, and ERNIE-Layout to assess the performance gap between multimodal LLMs and domain-specific fine-tuned models.
@@
The Baseline Configuration. The baseline configuration is outlined in Table 1, where the configuration is selected based on best practices, such as in Perot et al. (2024) for the following reasons: (1) OCR reflects real-world scenarios for digitized LRDs. (2) Medium chunk size balances efficiency and context preservation, addressing token limits in LLMs. (3) Few-shot prompting combines pre-trained knowledge with minimal task-specific guidance. (4) Using zero examples provides a clear benchmark for assessing the model’s raw performance. (5) Initial predictions are retained to evaluate models’ raw output without modifications, ensuring a direct assessment of their capabilities. (6) Finally, exact match provides a stringent measure of correctness, offering a reliable baseline for comparison across configurations.